OpenStudy (anonymous):

is this right? 12x^2+5x=3 -3 12x^2+5x-3=0 (3x+1)(4x-3) 3x+1=0 4x-3=0 -1 +3 3x=-1 4x=3 x=-1/3 x=3/4 if not tell me were i went wrong

OpenStudy (anonymous):

should be (3x-1)(4x+3)

OpenStudy (anonymous):

okay bc its a positive 5 right?

OpenStudy (anonymous):

You're right except your signs are backward.

OpenStudy (anonymous):

okay, i have a problem with my signs, i really get confused on were to place them

OpenStudy (anonymous):

\[(3x-1) = 0 \iff x = 1/3\] \[(4x + 3) = 0 \iff x = -3/4\] \[(3x-1)(4x+3) = 0 \iff x = 1/3\ \text{ or }\ x = -3/4\] And yeah, if you foil it out you'll see that you have 9x - 4x = 5x for the middle term.

OpenStudy (anonymous):

You can always check your factorization by foiling it out to see if you got what you started with. \[(3x+1)(4x-3) = 12x^2 -9x + 4x -3 = 12x^2 -5x -3\]

OpenStudy (anonymous):

Which isn't right.

OpenStudy (anonymous):

whats not right??

OpenStudy (anonymous):

(3x+1)(4x-3) isn't the factored form of \( 12x^2+5x−3\)